# Unsupervised Text Embedding Space Generation Using Generative Adversarial Networks for Text Synthesis

### v1

@@ 0.B.1 Optimization
The TESGAN framework performs adversarial training after training the seed interpretation model with multi-turn data. The multi-turn data for seed interpretation model training are limited to a maximum of four turns, and the maximum length of each sentence is 16. In the adversarial training phase, the sentences used as the seeds are composed of tokens explained in Equation 2, and the length of each sentence is set to 16 including special tokens. The discriminators are updated by the loss of SSD and SOD during adversarial training. Also, the generator is updated not only by the loss of SSD and SOD but also that of SDP and SFP.
Additionally, we find that the convergence of the model depends on the parameter update rate of the discriminators and the generator. According to Algorithm 1, the discriminators update parameters only at odd training epochs so that the generator can follow the learning of the discriminators; because the convergence of the generator is slower than that of the discriminator. If the discriminators are updated every training epoch in the same way as the generator, adversarial training is unbalanced.
Also, we conduct additional experiments by changing the number of the generator updates from one to three for each mini-batch step. If the generator updates only once per step as same as the discriminators, the generator cannot follow the learning of the discriminators. On the other hand, if the generator updates three times per step, the discriminators cannot follow the learning of the generator. Therefore, we update the generator twice per step. Consequently, the generator is updated four times more than the discriminators per two epochs. Algorithm 1 shows the entire training process of TESGAN.
0.B.2 Hyperparameters
The 12-layer GPT-2 is used as the seed interpretation model, and both hidden and embedded dimensions are 768. We adopt the byte-pair-encodings (BPE) [24] tokenizer with 50,260 vocabularies in the seed interpretation model. We suggest using the Adam optimizer [14] with 0.001 learning rate, and 100 mini-batch sizes for training the seed interpretation model.
@@ 0.B.2 Hyperparameters
The fake seeds are generated from the uniform distribution noise $X$ with an interval of $[-10,10)$ by the generator, which has two convolutional processes. In addition, the Leaky ReLU with slope 0.5 and the sigmoid are used in the middle and the end of the generator, respectively. We suggest the Adam optimizer with 0.0002 learning rate because the generator has difficulty converging when the learning rate exceeds 0.0004. The BERT and the LSTM models, used as SSD and SOD respectively, consist of two layers and their hidden dimension is 768. Both discriminators use the Adam optimizer with 0.0005 and 0.001 learning rate, respectively. When the learning rates of the discriminators are larger than the proposed values, adversarial learning is unbalanced. Also, a mini-batch size is set to 128 during adversarial training. Lastly, all the above experiments took place on a machine with Ubuntu 18.04.5 and an NVIDIA RTX 3090 GPU.

### v4

@@ Appendix D Hyperparameters
In the adversarial training phase, the sentences used as the seeds are composed of tokens explained in Equation 3, and the length of each sentence is set to 16 including special tokens. The discriminators are updated by the loss of SSD and SOD during adversarial training. Also, the generator is updated not only by the loss of SSD and SOD but also that of SDP and SFP.
@@
The fake seeds are generated from the uniform distribution noise $X$ with an interval of $[-10,10)$ by the generator, which has two convolutional processes. In addition, the Leaky ReLU with slope 0.5 and the sigmoid are used in the middle and the end of the generator, respectively. We used the Adam optimizer with $2e^{-4}$ learning rate when training DailyDialog because the generator has difficulty converging when the learning rate exceeds $4e^{-4}$. However, when training on IMDb, a larger dataset than DailyDialog, we set the learning rate to $5e^{-4}$. The BERT and the LSTM models, used as SSD and SOD respectively, consist of two layers and 768 hidden dimensions. Both discriminators used the Adam optimizer with $5e^{-4}$ and $1e^{-3}$ learning rate, respectively. When the learning rates of the discriminators were larger than the proposed values, adversarial learning was imbalanced. Also, the mini-batch size was set to 128 during adversarial training. Lastly, all the above experiments took place on a machine with Ubuntu 18.04.5 and an NVIDIA RTX 3090 GPU.
Appendix E Additional Results
We provide evaluation results of the text generated by each model per epoch. Table 8 shows the results of 1, 5, 10, 15, 17, and 20-epoch results of each model. Also, Figure 10 shows LM and SBL results of the TESGAN-based models and the baselines. In Figure 10, the SBL results of the baselines tend to increase.
Method Epoch FBD $\downarrow$ MSJ2 $\uparrow$ MSJ3 $\uparrow$ MSJ4 $\uparrow$ MSJ5 $\uparrow$ DSR ($R_{syn},R_{unq}$) $\uparrow$
TESGAN 1 3.441 0.129 0.067 0.034 0.018 0.911 (1, 0.836)
5 3.826 0.137 0.072 0.037 0.018 0.932 (1, 0.873)
10 3.185 0.132 0.069 0.035 0.016 0.932 (0.999, 0.873)
15 2.624 0.146 0.080 0.041 0.020 0.961 (1, 0.925)
17 2.899 0.150 0.079 0.042 0.021 0.967 (1, 0.936)
20 3.339 0.131 0.068 0.035 0.017 0.932 (1, 0.872)
P-TESGAN 1 6.146 0.112 0.057 0.029 0.013 0.803 (1, 0.671)
5 3.746 0.118 0.060 0.030 0.016 0.801 (0.998, 0.669)
10 2.274 0.131 0.066 0.032 0.014 0.841 (0.997, 0.727)
15 2.132 0.121 0.061 0.030 0.014 0.812 (1, 0.683)
17 2.274 0.122 0.063 0.031 0.014 0.789 (0.998, 0.653)
20 2.309 0.119 0.058 0.029 0.014 0.784 (0.997, 0.646)
SeqGAN 1 6.153 0.185 0.091 0.040 0.015 0.880 (0.883, 0.877)
5 6.373 0.121 0.065 0.032 0.014 0.602 (0.639, 0.568)
10 9.432 0.064 0.034 0.017 0.007 0.289 (0.34, 0.251)
15 18.471 0.012 0.006 0.003 0 0.019 (0.025, 0.015)
17 24.504 0.004 0.002 0.001 0 0.004 (0.006, 0.003)
20 28.048 0 0 0 0 0.001 (0.013, 0.001)
RankGAN 1 6.409 0.189 0.096 0.048 0.023 0.890 (0.895, 0.886)
5 6.778 0.160 0.084 0.04 0.016 0.82 (0.851, 0.791)
10 10.862 0.138 0.074 0.037 0.018 0.739 (0.799, 0.687)
15 9.732 0.118 0.060 0.030 0.015 0.699 (0.791, 0.625)
17 9.893 0.115 0.058 0.028 0.012 0.696 (0.792, 0.62)
20 12 0.114 0.056 0.026 0.011 0.721 (0.836, 0.634)
MaliGAN 1 21.436 0.015 0.006 0.003 0 0.030 (1, 0.015)
5 16.589 0.003 0 0 0 0.027 (1, 0.014)
10 57.769 0 0 0 0 0 (1, 0)
15 57.769 0 0 0 0 0 (1, 0)
17 57.769 0 0 0 0 0 (1, 0)
20 57.769 0 0 0 0 0 (1, 0)
PG-BLEU 1 9.002 0.071 0.036 0.015 0.006 0.569 (0.555, 0.584)
5 18.974 0.017 0.008 0 0 0.139 (0.472, 0.082)
10 21.568 0.001 0.001 0 0 0.041 (0.792, 0.021)
15 89.764 0 0 0 0 0.004 (0.773, 0.002)
17 142.384 0 0 0 0 0.003 (1, 0.001)
20 142.384 0 0 0 0 0.001 (1, 0.001)
Table 8: Performance of each model per epoch.
(a) LM $\downarrow$
(b) SBL3 $\downarrow$
(c) SBL4 $\downarrow$
Figure 10: LM, SBL results of TESGAN-based models and baselines trained with DailyDialog.